# Output Prediction of Quantum Circuits based on Graph Neural Networks

### v1

@@ 1 Introduction
Quantum computing, as a new computational paradigm, has the potential to address problems beyond the capabilities of classical computing with higher efficiency and speed. It has demonstrated exponential or polynomial advantages in various domains, including combinatorial optimization [14, 3, 20], large-scale communication systems [33, 17, 19, 18], molecular dynamics [29, 40], quantum chemistry[5, 34, 44], and machine learning [7, 11, 21, 23, 31, 30, 49]. Driven by breakthroughs in physical implementation technologies, quantum hardware has undergone rapid development over the past two decades, with numerous quantum computing systems now available [24, 1, 8, 2]. Despite its promising prospects, quantum computing must navigate an extended era of Noisy Intermediate-Scale Quantum (NISQ) devices before entering the fault-tolerant quantum computing era [41]. During this stage, the limitations imposed by coherence times and quantum gate errors pose significant bottlenecks to achieving quantum advantage. Quantum devices currently encounter limitations in resource accessibility, such as runtime and qubit count, which do not match the convenience provided by classical computers. Additionally, classical simulation methods often suffer from inefficiencies. Consequently, some advancements must proceed without adequate benchmark data, significantly impeding the engineering of quantum devices.
@@
Against this backdrop, machine learning presents an attractive alternative to direct classical simulation. For instance, in [48], Graph Transformers are utilized to predict parameters such as the Probability of Successful Trials (PST), which positively correlates with circuit fidelity, allowing for the early rejection of low-fidelity circuits. In [10], a meticulously designed qubit-scalable Convolutional Neural Network (CNN) is employed to predict expectation values of circuits, delivering results that, under specific conditions, outperform those obtained from the freely available Noisy Intermediate-Scale Quantum (NISQ) devices. In [28], neural networks based on simple Multilayer Perceptrons (MLPs) are used to predict kernel-target alignment (KTA), a cost-effective proxy for quantum kernel classification accuracy, facilitating the identification of high-accuracy quantum kernel circuits. Beyond these applications, estimating or predicting quantum circuit outputs as a replacement for classical simulation continues to have extensive use cases. For example, circuit output predictors could be integrated into diffusion models, such as those discussed in [4], to guide the generation of high-performance circuits. Alternatively, these predictors could support reinforcement learning (RL) processes, as outlined in [39], by evaluating reward functions, thereby reducing resource consumption during training.
Based on the considerations outlined above, we aim to predict the outputs of quantum circuits by designing a quantum circuit output predictor. Since the connections between quantum gates in a quantum circuit can be represented as directed acyclic graphs (DAGs), it is intuitive to employ Graph Neural Networks (GNNs) for output prediction. GNNs necessitate the construction of individual feature vectors for each node, which naturally facilitates the embedding of various attributes of quantum gates, such as gate type, gate error, and target qubits. This paper primarily investigates the application of GNNs for estimating or predicting the outputs of quantum circuits.
@@
We propose a GNN-based framework for predicting single-qubit and two-qubit expectation values of quantum circuits, as illustrated in figure 1. These limited expectation values are sufficient for circuits that require only one or two possible output bitstrings, such as the Bernstein–Vazirani (BV) algorithm [6] and the Deutsch–Jozsa algorithm [13]. Utilizing a parameter-free quantum gate set, we randomly generate quantum circuits under both noiseless and noisy conditions. Classical simulations are conducted using the Qiskit library [45] to obtain single-qubit and two-qubit expectation values for our dataset. During the transformation of random circuits into graphs, we design feature vectors for each node to encapsulate properties such as gate type, target qubits, and noise information. Our experiments demonstrate that the GNNs achieve strong predictive performance on datasets with a relatively small circuit search space. In comparison to convolutional neural networks (CNNs)-based methods, our GNNs-based method exhibits superior performance, greater flexibility in handling input circuit structures, and enhanced scalability.
@@
Next, we extend the GNNs-based prediction framework in two significant ways. First, we incorporate parameterized quantum gates, which broadens the applicability of GNNs, particularly to parameterized quantum circuits (PQCs). Second, the prediction target shifts to the overall property of a quantum circuit, specifically the ground state energy of a 4-qubit hydrogen molecule. We utilize the predicted ground state energy as a performance metric for randomly generated PQCs with various structures and propose two comparison schemes: Direct Comparison and Indirect Comparison of circuit performance. Simulation results under noisy conditions demonstrate that the Direct Comparison scheme significantly enhances prediction accuracy compared to the indirect approach.
@@ 3.1 GNNs for expectation values prediction
This is the data-driven approach, and intuitively, estimating expectation values does not require exactly computing the complete density matrix. Therefore, data-driven approaches present the opportunity to provide sufficiently accurate predictions at a lower computational cost. The framework for expectation prediction is divided into three primary stages. The first stage is to generate random quantum circuits with the assistance of a universal gate set and measure the circuit expectation values, which are used as training data for the GNNs. The second stage is to transform the generated random circuits into a graph structure and formulate feature vectors for each node that can contain device noise information. In the third stage, the graph is processed to predict circuit expectation values using GNNs.
This work has practical value as the framework realizes accurate prediction of the output expectation values of the circuits. When only a one-bit string has non-zero measurement probability in the circuit, the single-qubit expectation values are sufficient to find it. If combined with two-qubit expectation values, it is enough to identify the output bit strings when only two of them have non-zero probability [10]. Some relevant quantum algorithms, such as the Deutsch–Jozsa algorithm [13], the quantum counting algorithm [38], and the Grover algorithm with two searched items [15], can assist in solving problems related to the prediction of the expectation values mentioned above. In the following, we will discuss the specific implementation details of the framework outlined in figure 1.
@@ 3.1.1 Random circuits generation
To realize the first stage in our framework, random quantum circuits are generated to provide a database for subsequent GNNs training. We choose two single-qubit gates, namely the $\pi/8$ gate ($T$) and Hadamard gate ($H$), and one two-qubit gate CNOT, as a universal gate set to generate random circuits. The universal gate set consisting of parameter-free gates is denoted $S=\left\{T,H,CNOT\right\}$. In the first step, gates are randomly selected from a universal gate set $S$ to create an initial version of the random circuit with $N$ qubits and $P$ layers of gates, which we define as the circuit depth. The integer $N$ corresponds to the number of qubits and $P$ denotes the number of gates on each qubit. For the integrity of the circuits, we set the initial state of all qubits to $\left|0\right\rangle$, and the circuits include measurement operations. After completing the circuit construction, an optimization strategy is used to eliminate duplicate gates to obtain circuits for later training. These optimization strategies include the equivalence of two consecutive H-gates to an Identity gate, the equivalence of eight consecutive T-gates to an Identity gate, and the elimination of two consecutive CNOTs, etc. The specific process of generating random circuits is illustrated in figure 2.
@@ 3.1.3 Node features
We create a feature vector for each node in the graph to represent certain attributes of the node. The features include index, gate type, target qubit index, $T_{1}$ and $T_{2}$ of the target qubit, gate error, and readout error, as shown in figure 3.
@@
The length of the node feature vector is 31. The first part is the node index, which is used for node sorting and differentiation. The subsequent three parts are one-hot encoded representations of the node type, gate qubit, and tag. Here, the node type indicates the gate type: initial input qubit, measurement, CNOT, H, or T. For example, $00100$ signifies that the node represents a CNOT gate. The gate qubit is used to label the index of the qubit affected by the quantum gate, and its length can be expanded according to practical needs. The tag is effective when the gate is a CNOT, indicating the control qubit and the target qubit of the CNOT. One advantage of GNNs is their ability to consider hardware noise, primarily achieved through feature vectors. In this experiment, we utilize the last 7 numbers to characterize the noise properties of the hardware. The noise information is mainly referenced from IBM’s quantum devices, such as IBM Perth, IBM Lagos, IBM Nairobi, etc. The first 4 numbers respectively represent the $T_{1}$ and $T_{2}$ values for the first and second qubits. The 5th number denotes the gate error for the gate, while the last two numbers indicate readout errors. If a node does not possess a specific noise characteristic, it is set to 0. For example, for a measurement node, $T_{1}$, $T_{2}$, and gate error are all set to 0. For a node representing the CNOT, its complete feature vector is illustrated in figure 3.
@@ 3.2 GNNs for circuit performance comparison
In this chapter, we primarily expand upon the framework introduced in Section 3.1. Firstly, we replace the universal gate set for generating random quantum circuits, transitioning from the previous set $S=\left\{T,H,CNOT\right\}$ to $S^{\ast}=\left\{R_{x}\left(\theta\right),R_{y}\left(\beta\right),CNOT\right\}$. The updated universal gate set allows for the generation of parameterized quantum circuits (PQCs), significantly broadening the applicability of quantum circuits. For example, this tool can be utilized in the search for optimal PQC structures. Secondly, we extend the values predicted by the GNNs, enabling it not only to predict the expectation values of qubits (including single-qubit and two-qubit expectation values) but also to predict the overall properties of a quantum circuit. Finally, we choose the Variational Quantum Eigensolver (VQE) algorithm as an extension application of our Section 3.1 architecture. We achieve this by using randomly generated PQCs and the optimized parameters to determine the ground state energy values of the molecule, which are then used as data to train the GNNs. The ground state energy of the molecule is optimized through the following formula [43] by adjusting the parameters of different circuits:
@@ 4.1 Evaluation methodology
Dataset Setup. For the noisy and noiseless simulator datasets, we randomly generate 20000 random circuits for each case. In addition to noise information derived from real quantum devices (IBM Perch, IBM Lagos, IBM Nairobi, and IBM Jakarta), we also define a noise model, referred to as Simulated Noise, in which gate errors, $T_{1}$, $T_{1}$, and readout errors are fixed at predefined values. During the extrapolation process for qubits, we randomly generate 40000 random circuits as the dataset.
@@ 4.2.1 Single-qubit expectation values
$N$ Without Noise Simulated Noise IBM Perch IBM Lagos IBM Nairobi IBM Jakarta
@@ 4.2.3 Two-qubit expectation values
In addition to predicting single-qubit expectation values, GNNs can also be trained to predict two-qubit expectation values. We focus on the first two qubits, corresponding to the rescaled expectation value $z_{ij}$ defined in equation (7). Notably, the same GNNs model can predict the two-qubit expectation value for any pair of qubits by swapping the target qubits with the first two qubits. We randomly generated 20,000 circuits with depths ranging from 5 to 11 as the dataset and used GNNs to predict two-qubit expectation values under both noiseless conditions and noisy conditions with noise derived from IBM Perch. The results, including the coefficient of determination $R^{2}$, are presented in figure 12.

### v2

@@ 1 Introduction
Quantum computing, as a new computational paradigm, has the potential to address problems beyond the capabilities of classical computing with greater efficiency and higher speed. It has demonstrated exponential or polynomial advantages in various domains, including combinatorial optimization [1, 2, 3], large-scale communication systems [4, 5, 6, 7], molecular dynamics [8, 9], quantum chemistry[10, 11, 12], and machine learning [13, 14, 15, 16, 17, 18, 19]. Driven by breakthroughs in physical implementation technologies, quantum hardware has undergone rapid development over the past two decades, with numerous quantum computing systems now available [20, 21, 22, 23]. Despite its promising prospects, quantum computing must navigate an extended era of Noisy Intermediate-Scale Quantum (NISQ) devices before entering the fault-tolerant quantum computing era [24, 25]. During this stage, the limitations imposed by coherence times and quantum gate errors pose significant bottlenecks to achieving quantum advantage. Quantum devices currently encounter limitations in resource accessibility, such as runtime and qubit count, which do not match the convenience provided by classical computers. Additionally, classical simulation methods often suffer from inefficiencies. Consequently, some advancements must proceed without adequate benchmark data, significantly impeding the engineering of quantum devices. Against this backdrop, accurately predicting the output of quantum circuits is not only a theoretical challenge but also a crucial step toward optimizing quantum algorithms and enhancing hardware performance. However, due to the complex noise characteristics of quantum circuits and the computational challenges that grow with circuit size, existing classical simulation methods often struggle to provide efficient solutions.
@@
Notably, as it can provide low-cost predictions of specific aspects of quantum circuit outputs once the model is trained, machine learning has recently offered an attractive alternative to direct classical simulation. For instance, in [26], a meticulously designed qubit-scalable Convolutional Neural Network (CNN) is employed to predict expectation values of circuits, delivering results that, under specific conditions, outperform those obtained from the freely available Noisy Intermediate-Scale Quantum (NISQ) devices. In [27], neural networks based on simple Multilayer Perceptrons (MLPs) are used to predict kernel-target alignment (KTA), a cost-effective proxy for quantum kernel classification accuracy, facilitating the identification of high-accuracy quantum kernel circuits. Despite these advancements, CNN-based and MLP-based approaches face limitations in capturing the structural information of quantum circuits, particularly for large-scale systems or circuits with complex topologies. Graph-based models, on the other hand, naturally leverage the directed acyclic graph (DAG) representation of quantum circuits, making them a promising alternative for output prediction. For instance, in [28], Graph Transformers are utilized to predict the Probability of Successful Trials (PST), which positively correlates with circuit fidelity, achieving remarkable performance. Beyond these applications, estimating or predicting quantum circuit outputs still holds broad prospects. For example, circuit output predictors could be integrated into diffusion models, such as those discussed in [29], to guide the generation of high-performance circuits. Alternatively, these predictors could support reinforcement learning (RL) processes, as outlined in [30], by evaluating reward functions, thereby reducing resource consumption during training.
Based on the considerations outlined above, we aim to design a Graph Neural Networks (GNNs) predictor to predict the specific outputs of quantum circuits. This is not only because GNNs can capture the structural information of quantum circuits but also because they can construct individual feature vectors for each node, allowing the embedding of various properties of quantum gates, such as gate type, gate error, and target qubits. In this work, we primarily predict the single-qubit expectation values, two-qubit expectation values, and the overall property of quantum circuits.
@@
To begin with, we propose a GNN-based framework for predicting single-qubit and two-qubit expectation values of quantum circuits, as illustrated in figure 1. These limited expectation values are sufficient for circuits that require only one or two possible output bitstrings, such as the Bernstein–Vazirani (BV) algorithm [31] and the Deutsch–Jozsa algorithm [32]. Utilizing a parameter-free quantum gate set, we randomly generate quantum circuits under both noiseless and noisy conditions. Classical simulations are conducted using the Qiskit library [33] to obtain single-qubit and two-qubit expectation values for our dataset. During the transformation of random circuits into graphs, we design feature vectors for each node to encapsulate properties such as gate type, target qubits, and noise information. Our experiments demonstrate that the GNNs achieve strong predictive performance on datasets with a relatively small circuit search space. In comparison to convolutional neural networks (CNNs)-based methods, our GNNs-based method exhibits superior performance, greater flexibility in handling input circuit structures, and enhanced scalability.
@@
Furthermore, we extend our GNN-based prediction framework in two significant ways. First, we incorporate parameterized quantum gates, expanding its applicability to parameterized quantum circuits (PQCs). Second, instead of solely predicting expectation values, we shift our focus to evaluating the overall property of quantum circuits, specifically predicting the ground state energy of a 4-qubit hydrogen molecule. To compare the performance of PQCs, we propose two schemes: Direct Comparison, which directly predicts the relative performance difference between circuits, and Indirect Comparison, which involves predicting absolute ground state energies before comparison. Our simulation results demonstrate that the Direct Comparison scheme significantly outperforms the Indirect Comparison scheme, offering an average improvement of 36.2% on the same dataset. These findings highlight the potential of GNNs in efficiently evaluating PQC performance, providing a novel and effective perspective for analyzing quantum circuits.
@@ 3.1 GNNs for expectation values prediction
This is the data-driven approach, and intuitively, estimating expectation values does not require computing the complete density matrix exactly. Therefore, data-driven approaches present the opportunity to provide sufficiently accurate predictions at a lower computational cost. The framework for predicting expectation values is divided into three primary stages. The first stage is to generate random quantum circuits with the assistance of a universal gate set and measure the circuit expectation values, which are used as training data for the GNNs. The second stage is to transform the generated random circuits into a graph structure and formulate feature vectors for each node that can contain device noise information. In the third stage, the graph is processed to predict circuit expectation values using GNNs.
This work has practical value as the framework realizes accurate prediction of the output expectation values of the circuits. When only one-bit string has non-zero measurement probability in the circuit, the single-qubit expectation values are sufficient to find it. When combined with two-qubit expectation values, it is enough to identify the output bit strings when only two of them have non-zero probability [26]. Some relevant quantum algorithms, such as the Deutsch–Jozsa algorithm [32], the quantum counting algorithm [34], and the Grover algorithm with two searched items [51], can assist in solving problems related to the prediction of the expectation values mentioned above. In the following, we will discuss the specific implementation details of the framework outlined in figure 1.
@@ 3.1.1 Random circuits generation
To realize the first stage in our framework, random quantum circuits are generated to provide a database for subsequent GNNs training. We select two single-qubit gates, namely the $\pi/8$ gate ($T$) and Hadamard gate ($H$), and one two-qubit gate CNOT, as the universal gate set to generate random circuits. The universal gate set consisting of parameter-free gates is denoted $S=\left\{T,H,CNOT\right\}$. In the first step, gates are randomly selected from a universal gate set $S$ to create an initial version of the random circuit with $N$ qubits and $P$ layers of gates, which we define as the circuit depth. The integer $N$ corresponds to the number of qubits, and $P$ denotes the number of gates on each qubit. To ensure circuits integrity, we set the initial state of all qubits to $\left|0\right\rangle$, and the circuits include measurement operations. After completing the circuit construction, an optimization strategy is used to eliminate duplicate gates to obtain circuits for later training. These optimization strategies include the equivalence of two consecutive H-gates to an Identity gate, the equivalence of eight consecutive T-gates to an Identity gate, and the elimination of two consecutive CNOTs, etc. The specific process of generating random circuits is illustrated in figure 2.
@@ 3.1.3 Node features
We construct a feature vector for each node in the graph to represent certain attributes of the node. The features include index, node type, gate qubit, tag, the relaxation time ($T_{1}$) and dephasing time ($T_{2}$) of the target qubits, gate error, and readout error, as shown in figure 3.
@@
The length of the node feature vector is 31. The first part is the index of node, which is used for node sorting and differentiation. The subsequent three parts are one-hot encoded representations of the node type, gate qubit, and tag. Here, the node type indicates the gate type: initial input qubit, measurement, CNOT, H, or T. For example, $00100$ signifies that the node represents a CNOT gate. The gate qubit is used to label the index of the qubit affected by the quantum gate, and its length can be expanded according to practical needs. The tag is effective when the gate is a CNOT, indicating the control qubit and the target qubit of the CNOT. One advantage of GNNs is their ability to consider hardware noise, primarily achieved through feature vectors. In this experiment, we utilize the last 7 numbers to characterize the noise properties of the hardware. The noise information is mainly referenced from IBM’s quantum devices, such as IBM Perth, IBM Lagos, IBM Nairobi, etc. The first 4 numbers respectively represent the $T_{1}$ and $T_{2}$ values for the first and second qubits. The 5th number denotes the gate error for the gate, while the last two numbers indicate readout errors. If a node does not possess a specific noise characteristic, it is set to 0. For example, for a measurement node, $T_{1}$, $T_{2}$, and gate error are all set to 0. For a node representing the CNOT, its complete feature vector is illustrated in figure 3.
@@ 3.2 GNNs for circuit performance comparison
In this chapter, we primarily expand upon the framework introduced in Section 3.1. Firstly, we replace the universal gate set for generating random quantum circuits, transitioning from the previous set $S=\left\{T,H,CNOT\right\}$ to $S^{\ast}=\left\{R_{x}\left(\theta\right),R_{y}\left(\beta\right),CNOT\right\}$. The updated universal gate set allows for the generation of parameterized quantum circuits (PQCs), significantly broadening the applicability of quantum circuits. For example, this tool can be utilized in the search for optimal PQC structures. Secondly, we extend the values predicted by the GNNs, enabling them not only to predict the expectation values of qubits (including single-qubit and two-qubit expectation values) but also to predict the overall properties of a quantum circuit. Finally, we choose the Variational Quantum Eigensolver (VQE) algorithm as an extension application of our Section 3.1 architecture. We achieve this by using randomly generated PQCs and the optimized parameters to determine the ground state energy values of the molecule, which are then used as data to train the GNNs. The ground state energy of the molecule is optimized through the following formula [52] by adjusting the parameters of different circuits:
@@ 4.1 Evaluation methodology
Dataset Setup. For the noisy and noiseless simulator datasets, we randomly generate 20000 random circuits for each case. In addition to noise information derived from real quantum devices (IBM Perth, IBM Lagos, IBM Nairobi, and IBM Jakarta), we also define a noise model, referred to as Simulated Noise, in which gate errors, $T_{1}$, $T_{2}$, and readout errors are fixed at predefined values. During the extrapolation process for qubits, we randomly generate 40000 random circuits as the dataset.
@@ 4.2.1 Single-qubit expectation values
$N$ Without Noise Simulated Noise IBM Perth IBM Lagos IBM Nairobi IBM Jakarta
@@ 4.2.3 Two-qubit expectation values
In addition to predicting single-qubit expectation values, GNNs can also be trained to predict two-qubit expectation values. We focus on the first two qubits, corresponding to the rescaled expectation value $z_{ij}$ defined in equation (7). Notably, the same GNNs model can predict the two-qubit expectation value for any pair of qubits by swapping the target qubits with the first two qubits. We randomly generate 20,000 circuits with depths ranging from 5 to 11 as the dataset and used GNNs to predict two-qubit expectation values under both noiseless conditions and noisy conditions with noise derived from IBM Perth. The results, including the coefficient of determination $R^{2}$, are presented in figure 12.